Two crates, of mass m1= 50 kg and m2= 100 kg, are in contact and at rest on a horizontal surface. Force F= 600 N is exerted on the 50-kg crate. The coefficient of kinetic friction is 0.20. A) determine the acceleration of the system. B) determine the magnitude of the force that each crate exerts on the other.
Answer #1

Solution)

Given,

Mass, m1=50 kg

Mass, m2=100 kg

Force, F=600 N

Mass of crate=50 kg

Uk=0.20

Part a)

Consider Equilibrium of Forces

We have,

Force, F= uk(m1+m2)*g + a*(m1+m2)

Substituting values,

600 N=0.2(50+100)*9.8 +a*(50+100)

So, acceleration, a=2.04 m/s^2 (Ans)

=========

Part b)

Consider Mass m1 for Free Body Diagram,

We have, F'= F*m2/(m1+m2)

So, force exeeted on each other, F'=600*100/(50+100) =400 N (Ans)

===============